hannah is 2\5 of her mother age her mom is 35 years of age how old is hannah
Answer by Theo(13342)

h = hannah's age.
m = her mom's age.

formula is h = 2/5 * m

when m = 35, the formula becomes h = 2/5 * 35.

solve for h to get h = 2/5 * 35 = 2 * 35/5 = 2 * 7 = 14.